OpenStudy (lilai3):

Ms.Morris needs a student to help her with a demonstration for her class of 12 girls and 14 boys. She randomly chooses a student. What is the probability that she chooses a girl?

OpenStudy (anonymous):

12/14

OpenStudy (lilai3):

but i thought we had to add it al together....

OpenStudy (kinggeorge):

You do have to add it all together. So it would be 12/(12+14)=12/26=6/13

OpenStudy (lilai3):

thank you.. KingGeorge u were of great help and i thiank u for dat.... THANK YOU! and um.. i've gotta go study now.. for my probability test.. now.. so bye.. and good nite.. and thank you again! :D <333

OpenStudy (kinggeorge):

good night :) Feel free to tag me if you need my help later. If I'm around I'll help.

OpenStudy (lilai3):

thank you